# Beyond Wide-Angle Images: Unsupervised Video Portrait Correction via Spatiotemporal Diffusion Adaptation

### v1

@@ 3.3.1 Correction Trajectory
$R(t)=r(1)+r(2)+\cdots+r(t)$ (10)
where $r(1)$ is defined as a zero matrix. The rectified motion trajectory can be obtained by sequentially concatenating $R(1),R(2),\ldots,R(t),\ldots,R(N)$ over time. We achieve smooth trajectory stabilization and resolve video shaking issues by controlling the incremental stability of this pixel-level dense trajectory.
@@ 3.3.2 Spatiotemporal consistency
$\mathcal{L}_{video}=\mathcal{L}_{spatial}+\lambda\mathcal{L}_{temporal}$ (11)
For the temporal smoothness, we use the previously derived video warping motion trajectories to achieve this functionality, constraining the video trajectories obtained from the domain-adapted ImagePD model to be as smooth as possible. For three consecutive frames, we encourage the trajectory position of the middle frame to lie between the trajectory positions of the preceding and following frames:
$\mathcal{L}_{temporal}=\|R(t+1)+R(t-1)-2R(t)\|$ (12)
For spatial consistency, we leverage the pre-trained ImagePD model to ensure the quality of facial distortion correction within each frame. This is achieved by incorporating pseudo-labels generated by ImagePD, supervising the optical flow loss $\mathcal{L}_{flow}$ (Eq. 6) and the mask-based Sobel loss $\mathcal{L}_{mask}$(Eq. 8). These components maintain the structural integrity and edge details of the facial regions across individual frames, adapting the spatial correction capabilities from the image domain to the video domain.
@@ 3.4 Data preparation
Due to the lack of publicly available wide-angle video portrait datasets, we introduce a new dataset containing 136 clips, each with a resolution of 1080p at 30fps, ranging from 5 seconds to 90 seconds in duration(totaling 43,903 frames). The clips were captured using the iPhone 13 Pro, Gopro Hero 7, Samsung Galaxy S24, and Oneplus ace 2v, covering a wide variety of scenarios such as campuses, parks, shopping malls, classrooms, and other everyday settings. These clips encompass diverse scenarios, including content captured by self-media creators and film clips extracted from cinematic works. The number of people appearing in each scene varies from 1 to 6, reflecting diverse human dynamics. To enhance the dataset’s diversity, we’ve also included footage shot in special conditions like nighttime environments. To authentically replicate real-world recording conditions, we deliberately introduced artificial shaking to simulate the instability inherent in handheld video capture. This design choice poses significant challenges for stabilization algorithms. Fig. 4 presents some examples of our wide-angle video portrait dataset.
4 Experiments
4.1 Dataset and Implement Details
Our methodology builds upon two core datasets: the established wide-angle image dataset [33] and our newly captured video dataset.
Image Data:
In our experiments, we use the dataset from Tan et al.[33], which consists of 5133 training images and 129 test images captured with five wide-angle smartphones.
@@ 4.1 Dataset and Implement Details
Training Process of ImagePD:
Our framework employs a two-stage training protocol. In the first stage, we utilize the Adam optimizer[16] with an initial learning rate of $1\times 10^{-4}$, training the model for 200 epochs. The subsequent texture refinement stage leverages high-dimensional features extracted from the initial phase. This feature, combined with image $I_{s}$ and mask $M$, acts as the conditional guidance for the diffusion model to enhance fine-grained texture details. This stage implements a learning rate of $2\times 10^{-4}$, with optimization conducted over 500,000 iterations.
Training Process of VideoPD:
For VideoPD, we fine-tune the ImagePD model to maintain correction efficacy while smoothing temporal jitter. Training employs the Adam optimizer[16] with a learning rate of $2\times 10^{-4}$ over 300,000 iterations. The entire training process is conducted on a single RTX 4090.
4.2 Evaluation Metrics
Following Tan et al.’s work  [33], we evaluate geometric correction performance using two established metrics: LineAcc quantifies how well straight edges retain geometric integrity by computing deviations between local slopes and the ideal global orientation. ShapeAcc assesses shape preservation via cosine similarity between reference and corrected landmarks, where higher values indicate better alignment. We also employ Jin et al.’s Stability Score (Avg, Trans, Rot)[4], which uses homography and FFT, to measure VideoPD’s shake reduction, with higher scores denoting smoother videos.
4.3 Comparisons.
@@ 4.2 Evaluation Metrics
Shih et al.[29] TOG2019 66.143 97.253
Tan et al.[33] CVPR2021 66.784 97.490
Zhu et al.[47] CVPR2022 66.825 97.491
MOWA[19] arXiv 2024 - 97.475
DualPriors[42] ECCV2024 67.304 99.012
CoupledTPS[23] TPAMI2024 66.808 97.500
ImagePD (Ours) - 66.898 97.508
Table 1: Quantitative comparison of the proposed ImgaePD and different portrait correction methods.Higher LineAcc and ShapeAcc indicate better straight-line and face correction performance, respectively.
Figure 5: Comparison of ImagePD with existing methods on wide-angle distorted images. ImagePD achieves superior restoration of natural facial features and reduced edge distortions compared to state-of-the-art alternatives.
Image Comparisons. As shown in Table  1, our ImagePD achieves the second-best performance across LineAcc and ShapeAcc metrics, surpassed only by DualPriors[42]. This shows that our image correction model can retain the geometric integrity of straight edges and preserve facial shape well. Furthermore, in the qualitative comparison presented in Fig. 5, we can further observe that ImagePD outperforms DualPriors[42] in the aforementioned aspects. This undoubtedly demonstrates that, despite DualPriors[42] achieving state-of-the-art performance on certain metrics, their approach exhibits noticeable abrupt artifacts at stitching boundaries and inconsistent head-to-body proportions in the images. This may be attributed to their method relying on a parallel network that separately rectifies the background and portrait before stitching them into a single image, with blank areas repaired using LaMa[32]. In contrast, Fig. 5(c) clearly illustrates that our method generates more natural and refined character appearances, highlighting the superiority of ImagePD.
Video Comparisons. In the video domain, our approach not only corrects individual frames but also prioritizes inter-frame continuity to mitigate severe visual jitter. As demonstrated in Table 2, VideoPD exhibits superior performance in handling rotational distortions while maintaining optimal stability. In Fig. 6, the red arrows and boxes highlight the unnatural or failed corrections in previous methods when processing video sequences, and these errors typically result in video discontinuity and poor visual perception. In contrast, our framework effectively corrects distortions while preserving the spatiotemporal consistency across the entire video sequence, underscoring the superiority of VideoPD. Additional video results can be found in the supplementary materials.
@@
CoupledTPS[23] 0.9811 0.9785 0.9836
DualPriors[42] 0.9477 0.9290 0.9664
VideoPD(Ours) 0.9886 0.9821 0.9951
@@
Notably, the impressive performance of our proposed method ImagePD and VideoPD is primarily attributed to the two-stage structural design. This framework combines the transformer’s capacity for long-range dependency modeling with the diffusion model’s multi-step denoising process, achieving global structural consistency and local refinement. By integrating these complementary mechanisms within a unified architecture, our approach achieves remarkable generation quality while maintaining spatial-temporal consistency across various distortion scenarios.
4.4 Ablation studies
@@
Impact of Co-Designed Architecture and Conditional Guidance. Specifically, to independently analyze the impact of each component on correction quality, we designed four ablation scenarios: 1) Transformer Only: The Transformer predicts optical flow without diffusion, testing its global structure modeling. 2)Diffusion with Condition ($I_{s}$): The diffusion model exclusively uses the distorted image $I_{s}$ to predict optical flow, assessing its baseline detail correction. 3) Diffusion with Condition ($I_{s},M$): A mask is applied that marks the face region in the distorted wide-angle image, serving as the conditioning input to guide the model’s generation process. 4) Transformer + Diffusion with Condition ($I_{s},M$): The Transformer provides a feature vector $h$ concatenated with $I_{s}$ and $M$ to guide the diffusion model, evaluating their combined structural and detail refinement capabilities.
@@
Table 3: Ablations on the structure of co-designed architecture and different conditions.
Table  3 presents the experimental results, clearly demonstrating the impact of each module. Specifically, when the Transformer model is integrated with the diffusion model, the scores of LineAcc and ShapeAcc achieve their peak values. These findings confirm our hypothesis that the joint transformer-diffusion architecture can significantly improve the performance of wide-angle lens distortion correction. Furthermore, the progressive incorporation of conditioning information highlights how each component contributes to the robustness and accuracy of portrait stability.
@@
ImagePD 0.9325 66.898 97.508
VideoPD 0.9886 66.692 97.480
Table 4: Ablation Experiments Comparing ImagePD and VideoPD on Video Stabilization and Correction.
Impact of Spatiotemporal Smoothing on Video Correction. Our VideoPD model is a fine-tuned extension of the ImagePD model, specifically designed to address the challenges of video correction. Meanwhile, as mentioned before, the key difference between video and photo correction is the spatiotemporal consistency. Therefore, in this section, we directly compare the video results generated by VideoPD and ImagePD to evaluate whether VideoPD alleviates the warping shake caused by the correction process.
In fact, to mitigate the video shaking , some of the original correction of ImagePD will inevitably be compromised. In our experiments, we strive to balance correction and smoothing. This balance ensures that while we reduce the warping shake, we do not sacrifice too much of the correction accuracy. Tab. 4 presents the results. ImagePD achieves excellent correction accuracy, while VideoPD, with spatiotemporal smoothing, increases Stability to 0.9886 (a 6.0% improvement), significantly reducing jitter. The minor decrease in Line/ShapeACC is statistically insignificant ($p>0.05$, t-test), indicating that VideoPD effectively balances correction accuracy and inter-frame smoothness.
@@ 5 Conclusion
In this work, we introduce a novel approach to address facial distortion correction in both images and videos, marking the first application of diffusion models in the domain of portrait distortion correction. Our method represents the first deep learning-based solution for wide-angle video portrait correction, offering a significant advancement over traditional approaches. By leveraging image-level labels to learn the task of unlabeled wide-angle video correction, we drastically reduce the annotation burden and costs. Additionally, our proposed unsupervised framework, VideoPD, enhances spatiotemporal consistency while maintaining a smooth balance between correction and stabilization. We validate our method using a newly created wide-angle video dataset, demonstrating superior performance compared to existing approaches, showcasing its potential for wide-scale, automated, and accurate correction of wide-angle portrait distortion in still images and dynamic video sequences.

### v2

@@ 3.3 VideoPC
$R(t)=r(1)+r(2)+\cdots+r(t),$ (9)
where $r(1)$ is defined as a zero matrix. The final correction trajectory can be obtained by sequentially chaining $R(1),R(2),\ldots,R(t),\ldots,R(N)$ over time.
The produced trajectory contains comprehensive motions, including the shakes caused by unsmooth rectification flows and the inherent jitters in original wide-angle videos.
@@
$\mathcal{L}_{video}=\mathcal{L}_{spatial}+\lambda\mathcal{L}_{temporal}.$ (10)
For the temporal smoothness, we encourage the correction trajectory from VideoPC to be as smooth as possible. In particular, for three consecutive frames, we require the trajectory position of the middle frame to lie between the trajectory positions of the preceding and following frames:
$\mathcal{L}_{temporal}=\|R(t+1)+R(t-1)-2R(t)\|.$ (11)
For spatial consistency, we leverage the pre-trained ImagePC model to ensure the quality of facial distortion correction within each frame. Concretely, we use pseudo-labels generated by ImagePC to provide spatial supervision for VideoPC, including the optical flow constraint $\mathcal{L}_{flow}$ (Eq. 4) and mask-based edge constraint $\mathcal{L}_{mask}$(Eq. 6). In the transfer process of facial correction capabilities from the image to video domain, these components prevent the structural integrity and edge details of the facial regions from severe degradation.
@@ 3.4 Data preparation
Due to the lack of publicly available wide-angle video portrait datasets, we introduce a new dataset containing 136 clips, each with a resolution of 1080p at 30fps, ranging from 5 seconds to 90 seconds in duration (totaling 43,903 frames). Our dataset includes clips captured by seven smartphones (e.g., iPhone 13 Pro) covering a wide variety of scenarios and encompass diverse scenarios, including content captured by self-media creators and film clips extracted from cinematic works. Detailed information is available in the supplementary materials. To authentically replicate real-world recording conditions, we deliberately introduced artificial shaking to simulate the instability inherent in handheld video capture. This design choice poses significant challenges for stabilization algorithms. Fig. 4 presents some examples of our wide-angle video portrait dataset.
4 Experiment
4.1 Dataset and Implementation Detail
Our methodology builds upon two core datasets: the established wide-angle image dataset (Tan et al. 2021) and our newly captured video dataset.
Image Data
In our experiments, we use the dataset from Tan et al.(Tan et al. 2021), which consists of 5133 training images and 129 test images captured with five wide-angle smartphones.
@@ 4.1 Dataset and Implementation Detail
Figure 5: Comparison of ImagePC with existing methods on wide-angle distorted images. ImagePC achieves superior restoration of natural facial features and reduced edge distortions compared to state-of-the-art alternatives.
Training Process of ImagePC
Our framework employs a two-stage training protocol. In the first stage, we utilize the Adam optimizer(Kingma and Ba 2017) with an initial learning rate of $1\times 10^{-4}$, training the model for 200 epochs. The subsequent texture refinement stage leverages high-dimensional features extracted from the initial phase. This feature, combined with image $I_{s}$ and mask $M$, acts as the conditional guidance for the diffusion model to enhance fine-grained texture details. This stage implements a learning rate of $2\times 10^{-4}$, with optimization conducted over 500,000 iterations.
Training Process of VideoPC
For VideoPC, we fine-tune the ImagePC model to maintain correction efficacy while smoothing temporal jitter. Training employs the Adam optimizer(Kingma and Ba 2017) with a learning rate of $2\times 10^{-4}$ over 300,000 iterations. The entire training process is conducted on a single RTX 4090.
4.2 Evaluation Metric
Following Tan et al.’s work  (Tan et al. 2021), we evaluate geometric correction performance using two established metrics: LineAcc quantifies how well straight edges retain geometric integrity by computing deviations between local slopes and the ideal global orientation. ShapeAcc assesses shape preservation via cosine similarity between reference and corrected landmarks, where higher values indicate better alignment. We also employ Jin et al.’s Stability Score (Avg, Trans, Rot)(Choi and Kweon 2020), which uses homography and FFT, to measure VideoPC’s shake reduction, with higher scores denoting smoother videos.
4.3 Comparative Result
@@
Shih et al. TOG2019 66.143 97.253
Tan et al. CVPR2021 66.784 97.490
Zhu et al. CVPR2022 66.825 97.491
DualPriors ECCV2024 67.304 99.012
CoupledTPS TPAMI2024 66.808 97.500
MOWA TPAMI2025 - 97.475
ImagePC (Ours) - 66.898 97.508
Table 1: Quantitative comparison of the proposed ImgaePC with other portrait correction methods. Higher LineAcc and ShapeAcc indicate better straight-line and face correction performance, respectively.
Image Comparison As shown in Tab.  1, our ImagePC achieves the second-best performance across LineAcc and ShapeAcc metrics, surpassed only by DualPriors(Yao et al. 2024). This shows that our image correction model can retain the geometric integrity of straight edges and preserve facial shape well. Furthermore, in the qualitative comparison presented in Fig. 5, we can further observe that ImagePC outperforms DualPriors(Yao et al. 2024) in the aforementioned aspects. This undoubtedly demonstrates that, despite DualPriors(Yao et al. 2024) achieving state-of-the-art performance on certain metrics, their approach exhibits noticeable abrupt artifacts at stitching boundaries and inconsistent head-to-body proportions in the images. This may be attributed to their method relying on a parallel network that separately rectifies the background and portrait before stitching them into a single image, with blank areas repaired using LaMa(Suvorov et al. 2021). In contrast, Fig. 5(c) clearly illustrates that our method generates more natural and refined character appearances, highlighting the superiority of ImagePC.
Video Comparison In the video domain, our approach not only corrects individual frames but also prioritizes inter-frame continuity to mitigate severe visual jitter. As demonstrated in Table 2, VideoPC exhibits superior performance in handling rotational distortions while maintaining optimal stability. In Fig. 6 (a), the red arrows and boxes highlight the unnatural or failed corrections in previous methods when processing video sequences, and these errors typically result in video discontinuity and poor visual perception. In contrast, our framework effectively corrects distortions while preserving the spatiotemporal consistency across the entire video sequence, underscoring the superiority of VideoPC. Additional video results can be found in the supplementary materials.
Video Correction Trajectory In Fig. 6 (b), we arbitrarily select a pixel from video frames and plot its trajectory, revealing that face distortion correction introduces noticeable jitter. Our smoothing process effectively mitigates this warping shake, producing smoother and more stable trajectories compared to CoupledTPS (Nie et al. 2024a) and DualPriors+LAMA (Yao et al. 2024; Suvorov et al. 2021), as shown in an enlarged view. This results in corrected video frames that are more faithful to the original, significantly reducing jitter and enhancing the viewing experience.
@@
CoupledTPS 0.9811 0.9785 0.9836
DualPriors 0.9477 0.9290 0.9664
VideoPC(Ours) 0.9886 0.9821 0.9951
@@
Notably, the impressive performance of our proposed method ImagePC and VideoPC is primarily attributed to the two-stage structural design. This framework combines the transformer’s capacity for long-range dependency modeling with the diffusion model’s multi-step denoising process, achieving global structural consistency and local refinement. By integrating these complementary mechanisms within a unified architecture, our approach achieves remarkable generation quality while maintaining spatial-temporal consistency across various distortion scenarios.
4.4 Ablation Study
@@
Structure-to-Detail Model and Conditional Guidance: Specifically, to independently analyze the impact of each component on correction quality, we designed four ablation scenarios: 1) Transformer Only: The Transformer predicts optical flow without diffusion, testing its global structure modeling. 2) Diffusion with Condition ($I_{s}$): The diffusion model exclusively uses the distorted image $I_{s}$ to predict optical flow, assessing its baseline detail correction. 3) Diffusion with Condition ($I_{s},M$): A mask is applied that marks the face region in the distorted wide-angle image, serving as the conditioning input to guide the model’s generation process. 4) Transformer + Diffusion with Condition ($I_{s},M$): The Transformer provides a feature vector $h$ concatenated with $I_{s}$ and $M$ to guide the diffusion model, evaluating their combined structural and detail refinement capabilities.
@@
Table 3: Ablation experiments on the hybrid architecture and different conditions.
Tab.  3 presents the experimental results, clearly demonstrating the impact of each module. Specifically, when the Transformer model is integrated with the diffusion model, the scores of LineAcc and ShapeAcc achieve their peak values. These findings confirm our hypothesis that the joint transformer-diffusion architecture can significantly improve the performance of wide-angle lens distortion correction. Furthermore, the progressive incorporation of conditioning information highlights how each component contributes to the robustness and accuracy of portrait stability.
@@
ImagePC 0.9325 66.898 97.508
VideoPC 0.9886 66.692 97.480
Table 4: Ablation experiments comparing ImagePC and VideoPC on video stabilization and correction.
Spatiotemporal Smoothing on Video Correction: Our VideoPC model is a fine-tuned extension of the ImagePC model, specifically designed to address the challenges of video correction. Meanwhile, as mentioned before, the key difference between video and photo correction is the spatiotemporal consistency. Therefore, in this section, we directly compare the video results generated by VideoPC and ImagePC to evaluate whether VideoPC alleviates the warping shake caused by the correction process.
In fact, to mitigate the video shaking , some of the original correction of ImagePC will inevitably be compromised. In our experiments, we strive to balance correction and smoothing. This balance ensures that while we reduce the warping shake, we do not sacrifice too much of the correction accuracy. Tab. 4 presents the results. ImagePC achieves excellent correction accuracy, while VideoPC, with spatiotemporal smoothing, increases Stability to 0.9886 (a 6.0% improvement), significantly reducing jitter. The minor decrease in Line/ShapeAcc is statistically insignificant ($p>0.05$, t-test), indicating that VideoPC effectively balances correction accuracy and inter-frame smoothness.
@@ 5 Conclusion
In this work, we introduce a novel approach to address facial distortion correction in both images and videos, integrating of structure and detail for effective portrait distortion correction. Our method represents the first deep learning-based solution for wide-angle video portrait correction, offering a significant advancement over traditional approaches. By using image-level labels to learn the task of unlabeled wide-angle video correction, we drastically reduce the annotation burden and costs. Additionally, our proposed unsupervised framework, VideoPC, enhances spatiotemporal consistency while maintaining a smooth balance between correction and stabilization. We validate our method using a newly created wide-angle video dataset, demonstrating superior performance compared to existing approaches, showcasing its potential for wide-scale, automated, and accurate correction of wide-angle portrait distortion in still images and dynamic video sequences.